# Learning Distributions on Manifolds with Free-form Flows

### v1

@@ 4 Manifold free-form flows
The free-form flow (FFF) framework allows training any pair of parameterized encoder $f_{\theta}(x)$ and decoder $g_{\phi}(z)$ as a generative model [11], see Sec. 3.1. In this section, we first show how we set up our models to learn encoders and decoders that are restricted to the manifold. Then, we proceed to show how we adapt the FFF training to manifold free-form flows (M-FFF).
4.1 Learning a function on the manifold
For both encoder $f_{\theta}$ and decoder $g_{\phi}$, we employ feed-forward neural networks $\tilde{f}_{\theta},\tilde{g}_{\phi}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ that act on the embedding space $\mathbb{R}^{m}$ of the manifold $\mathcal{M}$. Since the data and latent codes both lie on the manifold, we project the outputs of the networks to the closest point on the manifold:
$f_{\theta}(x)=\operatorname{proj}(\tilde{f}_{\theta}(x)),$ (9)
$g_{\phi}(z)=\operatorname{proj}(\tilde{g}_{\phi}(z)).$ (10)
By restricting the input of both networks to the manifold, we obtain functions acting on the manifold:
$f_{\theta}(x):\mathcal{M}\to\mathcal{M},$ (11)
$g_{\phi}(z):\mathcal{M}\to\mathcal{M}.$ (12)
Figure 1 shows an example that maps from a latent uniform distribution to a bimodal distribution on the circle. In the next section, we explain how to estimate the gradient of the negative log-likelihood.
4.2 Manifold gradient estimator
The adaptation of free-form flows to isometrically embedded manifolds is remarkably simple. Due to the similarity of the change of variables formula (Eq. 8) to the Euclidean case, we can simply adapt the FFF estimator:
@@ 4.2 Manifold gradient estimator
$\nabla_{\theta}\mathcal{L}_{\text{NLL}}=\nabla_{\theta}\mathbb{E}_{x,v}\left[-\log p_{Z}(z)-v^{T}f_{\theta}^{\prime}(x)\texttt{SG}[g_{\phi}^{\prime}(z)v]\right]$ (13)
@@
where the only change is that $v$ must be sampled from the tangent space at $f_{\theta}(x)$, and must have unit covariance in that space. We achieve this by sampling standard normal vectors $\tilde{v}\in\mathbb{R}^{m}$ and then projecting them into the tangent space using the Jacobian of the projection function:
$v=\operatorname{proj}^{\prime}(f_{\theta}(x))\tilde{v}.$ (14)
We derive this estimator in detail in Sec. 7.2.
An equivalent formulation is to prepend $\tilde{f}$ and $\tilde{g}$ by a projection in addition to appending the projection:
$f_{\theta}=\operatorname{proj}\circ\tilde{f}_{\theta}\circ\operatorname{proj}$ (15)
and
$g_{\phi}=\operatorname{proj}\circ\tilde{g}_{\phi}\circ\operatorname{proj}.$ (16)
In this case $v$ can be directly sampled in $\mathbb{R}^{m}$, so that the Euclidean gradient estimator from Eq. 5 applies without modification. Nevertheless, we choose to use the first formulation ($v$ projected to the tangent space) as it allows us to reduce the variance of the estimate by normalizing $v$ after projection. A discussion of this point can found in Sec. 7.3.
4.3 Reconstruction loss and other regularization
@@
The reconstruction loss (Eq. 3) can be modified to use an on-manifold distance (e.g. great circle distance for the sphere) but we find that ambient Euclidean distance works well in practice, since it is almost identical for small distances. Hence we use the same Euclidean reconstruction loss as in free-form flows.
@@
$\mathcal{L}_{\text{U}}=\mathbb{E}_{x\sim\mathcal{U}(\mathcal{M})}[\|g_{\phi}(f_{\theta}(x))-x\|^{2}]$ (17)
which helps ensure that we have a globally consistent mapping between the data and latent manifolds in low data regions. Secondly, the squared distance between the output of $\tilde{f}_{\theta}$ and its projection to the manifold:
$\mathcal{L}_{\text{P}}=\mathbb{E}_{q(x)}[\|\tilde{f}_{\theta}(x)-f_{\theta}(x)\|^{2}]$ (18)
which discourages the output of $\tilde{f}_{\theta}$ from entering unprojectable regions, for example the origin when the manifold is $\mathbb{S}^{n}$. The same regularizations can be applied starting from the latent space.
@@
$\mathcal{L}=\mathcal{L}_{\text{NLL}}+\beta_{\text{R}}\mathcal{L}_{\text{R}}+\beta_{\text{U}}\mathcal{L}_{\text{U}}+\beta_{\text{P}}\mathcal{L}_{\text{P}}$ (19)
where $\beta_{\text{R}}$, $\beta_{\text{U}}$ and $\beta_{\text{P}}$ are hyperparameters. We give our choices in Sec. 8.
@@ 5 Experiments
In this section, we demonstrate the practical performance of manifold free-form flows on a variety of manifolds. We choose established experiments to ensure comparability with previous methods. Generally, we find that we obtain competitive performance with the advantage that our method samples orders of magnitudes faster because it requires only a single function evaluation. In contrast, previous general purpose work integrates a differential equation on the manifold, which requires multiple model evaluations.
We give all details on training in the appendix.
5.1 Synthetic distribution over rotation matrices
The group of 3D rotations $SO(3)$ can be represented by rotations matrices with positive determinant, i.e. all $Q\in\mathbb{R}^{3\times 3}$ with $Q^{T}Q=I$ and $\det Q=1$. To apply manifold free-form flows, we project an output matrix $R\in\mathbb{R}^{3\times 3}$ from the encoder/decoder to the subspace of special orthogonal matrices by finding the matrix $Q\in SO(3)$ minimizing the Frobenius norm $\|Q-R\|_{F}$. This is known as the constrained Procrustes problem and the solution $Q$ can be determined via the SVD $R=U\Sigma V^{T}$ [26]:
$Q=USV^{T},$ (20)
We learn a M-FFF on synthetic mixture distributions proposed by [10] with $M$ mixture components for $M=16,32$ and $64$. Samples from one of the distributions and samples from our model are depicted in Fig. 3.
M-FFF achieves negative log-likelihoods on par or better than previous methods, see Tab. 3. At the same time, it speeds up sampling by about two orders of magnitude since it requires only a single function evaluation to sample.
Figure 3: Manifold free-form flows on a synthetic $SO(3)$ mixture distribution with $M=64$ mixture components proposed by [10]. (Top) 10,000 samples each from the ground truth distribution and (bottom) our model. This visualization computes three Euler angles, which fully describe a rotation matrix, and then plot the first two angles on the projection of a sphere and the last by color [34]. We find that our model nicely samples from the distribution with few outliers between the modes.
$M=16$ $M=32$ $M=64$ Sampling steps
Moser flow [45] -0.85
@@ 5.1 Synthetic distribution over rotation matrices
$\sim$ 100
M-FFF (ours) -0.87
$\pm$ 0.02
-0.21 $\pm$ 0.02 0.45 $\pm$ 0.02 1
Table 3: Results on synthetic $SO(3)$ mixtures. We compare the final model’s negative log-likelihood (NLL, $\downarrow$) and the number of function evaluations necessary to obtain a single sample averaged over five runs $\pm$ standard deviations. Our manifold free-form flows (M-FFF) achieve comparable likelihoods at greatly improved sampling speed. Baselines values due to [10].
5.2 Earth data on the sphere
Volcano Earthquake Flood Fire Sampling steps
Mixture of Kent [41] -0.80
@@ 5.2 Earth data on the sphere
N/A
Riemannian CNF [30] -6.05
$\pm$ 0.61
0.14
$\pm$ 0.23
1.11
$\pm$ 0.19
-0.80
$\pm$ 0.54
$\sim$ 100
Moser flow [45] -4.21
$\pm$ 0.17
-0.16
$\pm$ 0.06
0.57
$\pm$ 0.10
-1.28
$\pm$ 0.05
$\sim 100$
Stereographic score-based [10] -3.80
$\pm$ 0.27
-0.19
$\pm$ 0.05
0.59
$\pm$ 0.07
-1.28
$\pm$ 0.12
$\sim$ 100
Riemannian score-based [10] -4.92
$\pm$ 0.25
-0.19
$\pm$ 0.07
0.45
$\pm$ 0.17
-1.33
$\pm$ 0.06
$\sim$ 100
Riemannian diffusion [20] -6.61
$\pm$ 0.97
-0.40 $\pm$ 0.05 0.43
$\pm$ 0.07
-1.38
$\pm$ 0.05
$>$ 100
Riemannian flow matching [8] -7.93 $\pm$1.67 -0.28
$\pm$ 0.08
0.42 $\pm$ 0.05 -1.86 $\pm$ 0.11 not reported
M-FFF (ours) -2.25
$\pm$ 0.02
-0.23
$\pm$ 0.01
0.51
$\pm$ 0.01
-1.19
$\pm$ 0.03
1
Table 4: Negative log-likelihoods for the earth datasets (lower is better). Results for the baselines taken from [10], [20] and [8]. While our results show mixed performance compared to the baselines, we uniquely provide a competitive single-step sampling method.
We evaluate manifold free-form flows on spheres with datasets from the domain of earth sciences. We use four established datasets compiled by [30] for density estimation on $\mathbb{S}^{2}$: Volcanic eruptions [38], earthquakes [37], floods [4] and wildfires [12].
We compare M-FFF to Riemannian score-based generative models [10], Riemannian diffusion models [20], Riemannian flow matching [8] and the other baseline models used in [10]. We provide a full comparison of negative log-likelihoods in Table 4 and visualize the learned densities in Figure 4. Our method has a mixed performance, but provides the only competitive solution that allows for single-step sampling. While it is outperformed by Riemannian flow matching in all datasets, the computational cost of training and evaluation is significantly smaller due to the single step sampling and greatly reduced model size (around 100 times smaller).
We notice that, since there are many regions without training data in the different datasets, the predicted latent codes and thus likelihood estimates of the validation set are often inaccurate. To ensure an accurate estimation of the data likelihood under the decoder, during evaluation we sample multiple latent points $z\sim\mathcal{N}(f_{\theta}(x),10^{-4})$ from a small region around the predicted latent point and project them back onto the manifold. To estimate the likelihood of $x$ we pick the point $z^{\star}=\operatorname*{arg\,min}_{z}d(x,g_{\phi}(z))$. This results in a reduction of the reconstruction error from $80$km to less than $10$km on average. We find that the likelihoods computed in this way now agree with the decoder distribution. Note that this does not affect sampling speed.
5.3 Torsion angles of molecules on tori
General Glycine Proline Pre-Pro RNA Sampling steps
Mixture of power spherical [20] 1.15
@@ 5.3 Torsion angles of molecules on tori
N/A
Riemannian diffusion models [20] 1.04
$\pm$ 0.012
1.97
$\pm$ 0.012
0.12
$\pm$ 0.011
@@
$\pm$ 0.004
-3.70
$\pm$ 0.592
$\sim$ 1000
Riemannian flow matching [8] 1.01
$\pm$ 0.025
1.90
$\pm$ 0.055
0.15
$\pm$ 0.027
1.18
$\pm$ 0.055
-5.20
$\pm$ 0.067
not reported
M-FFF (ours) 1.03
$\pm$ 0.0002
1.92
$\pm$ 0.0013
0.21
$\pm$ 0.0019
1.27
$\pm$ 0.004
-4.27
$\pm$ 0.09
1
Table 5: Negative log-likelihoods (NLL, $\downarrow$) on torus datasets. Experimental data for a mixture of 4096 power spherical distributions [7] and Riemannian diffusion models are taken from [20]. Furthermore, we compare against Riemannian flow matching [8]. Mean and standard deviation are evaluated over five independent models for each dataset. Notably, our method is able to achieve on par negative log-likelihoods with a single sampling step on most of the datasets and even outperforms Riemannian diffusion models on some of them.
Figure 5: Log density of M-FFF models in the $(\Phi,\Psi)$-plane of protein backbone dihedral angles. The test data for each dataset is shown as black dots. This representation was introduced by [42] and is known as Ramachandran plot. Note how the learned distribution adheres to the periodic boundary conditions.
@@
To benchmark manifold free-form flows on tori $\mathbb{T}^{n}$, we follow [20] and evaluate our model on two datasets from structural biology. We consider the torsion (dihedral) angles of the backbone of protein and RNA substructures respectively.
@@
We represent a tuple of angles $(\phi_{1},\dots,\phi_{n})\in[0,2\pi]^{n}$ by mapping each angle to a position on a circle: $X_{i}=(\cos\phi_{i},\sin\phi_{i})\in\mathbb{S}^{1}$. Then we stack all $X_{i}$ into a matrix $X\in\mathbb{R}^{n\times 2}$, compare Tab. 2.
@@
The first dataset is comprised of 500 proteins assembled by [29] and is located on $\mathbb{T}^{2}$. The three dimensional arrangement of a protein backbone can be described by the so called Ramachandran angles [42] $\Phi$ and $\Psi$, which represent the torsion of the backbone around the $N-C_{\alpha}$ and $C_{\alpha}-C$ bonds. The data is split into four distinct subsets General, Glycine, Proline and Pre-Proline, depending on the residue of each substructure.
@@
The second dataset is extracted from a subset of RNA structures introduced by [35]. As the RNA backbone structure can be characterized by seven torsion angles, in this case we are dealing with data on $\mathbb{T}^{7}$.
We record the negative log-likelihoods of our method in Tab. 5 and compare it to previous approaches. Notably, our approach demonstrates comparable performance across all datasets at fast sampling speed. In addition to the quantitative results, in Fig. 5 we show the log densities of the M-FFF models for the four protein datasets.

### v2

@@ 4 Manifold free-form flows
The free-form flow (FFF) framework allows training any pair of parameterized encoder $f_{\theta}(x)$ and decoder $g_{\phi}(z)$ as a generative model, see Section 3.1. In this section, we demonstrate how to generalize the steps in Section 3.1 to arbitrary Riemannian manifolds. Note that for simplicity, we choose the same manifold in data and latent spaces, i.e. $\mathcal{M}_{X}=\mathcal{M}_{Z}=\mathcal{M}$, but the method readily applies to $\mathcal{M}_{X}\neq\mathcal{M}_{Z}$ or $G_{X}\neq G_{Z}$ as long as they are topologically compatible, like a sphere and a closed 3D surface without holes. The detailed derivations in the appendix consider this generalization.
Manifold change of variables
The volume change on manifolds generalizes the Euclidean variant in Eq. 1 by (a) considering the change of volume in the tangent space and (b) accounting for volume change due to changes in the metric:
Theorem 2 (Manifold change of variables).
Let $(\mathcal{M},G)$ be a $n$-dimensional Riemannian manifold embedded in $\mathbb{R}^{m}$, i.e., $\mathcal{M}\subseteq\mathbb{R}^{m}$. Let $p_{X}$ be a probability distribution on $\mathcal{M}$ and let $f:\mathcal{M}\rightarrow\mathcal{M}$ be a diffeomorphism. Let $p_{Z}$ be the pushforward of $p_{X}$ under $f$ (i.e., if $p_{X}$ is the probability density of $X$, then $p_{Z}$ is the probability density of $f(X)$).
Let $x\in\mathcal{M}$. Define $Q\in\mathbb{R}^{m\times n}$ as an orthonormal basis for $\mathcal{T}_{x}\mathcal{M}$ and $R\in\mathbb{R}^{m\times n}$ as an orthonormal basis for $\mathcal{T}_{f(x)}\mathcal{M}$.
Then, the probability densities $p_{X}$ and $p_{Z}$ are related under the change of variables $x\mapsto f(x)$ by the following equation:
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|R^{T}f^{\prime}(x)Q|+\tfrac{1}{2}\log\frac{|R^{T}G(f(x))R|}{|Q^{T}G(x)Q|}.$ (9)
where $Q$ and $R$ depend on $x$ and $f(x)$, respectively, although this dependency is omitted for brevity.
Figure 2: Computation of the volume change in the tangent space of the manifold: The manifold change of variables formula in Eq. 10 requires to compute the change of a volume element in the tangent spaces under f, which in this example is given by the ratio of lengths of $dt$ and $dt^{\prime}$. Since $f$ is a map in the embedding space, $f^{\prime}(x)$ defines a linear map between vectors from the embedding space. To correctly compute the change in volume, we use $Q$ and $R$ to change coordinates to the intrinsic tangent spaces, resulting in the linear map $R^{T}f^{\prime}(x)Q:\mathcal{T}_{x}\mathcal{M}\rightarrow\mathcal{T}_{f(x)}\mathcal{M}$, which maps $dt$ to $dt^{\prime}$.
To give an intuition for this result, Fig. 2 shows how the volume change is computed for an isometrically embedded manifold, that is $G=I$ so that $|R^{T}GR|=|Q^{T}GQ|=1$. This simplifies Eq. 9 to:
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|R^{T}f^{\prime}(x)Q|.$ (10)
This is very similar to the familiar change of variables formula in the Euclidean case in Eq. 1, the only difference being that the determinant is evaluated on the $n\times n$ projection of $f^{\prime}(x)$ into the tangent spaces. These projections are necessary as the Jacobian of $f$ is singular in the embedding space, since its action is restricted to the local tangent spaces. See the full proof in Section A.1.
Manifold gradient estimator
We now generalize the volume change gradient estimator in Theorem 1 to an invertible function on the manifold $f_{\theta}:\mathcal{M}\to\mathcal{M}$. We find that taking the gradient of the manifold change of variables in Eq. 9 results in essentially the same computation as in the Euclidean case, but the trace in is now evaluated in the local tangent space:
Theorem 3.
Under the assumptions of Theorem 2 with $f=f_{\theta}$. Let $v\in\mathbb{R}^{m}$ be a random variable with zero mean and covariance $RR^{T}$. Then, the derivative of the change of variables term has the following trace expression, where $z=f_{\theta}(x)$:
$\nabla_{\theta}\log|R^{T}f_{\theta}^{\prime}(x)Q|$ $=\operatorname{tr}(R^{T}(\nabla_{\theta}f_{\theta}^{\prime}(x)){f_{\theta}^{-1}}^{\prime}(z)R)$ (11)
$=\mathbb{E}_{v}\left[v^{T}(\nabla_{\theta}f_{\theta}^{\prime}(x)){f_{\theta}^{-1}}^{\prime}(z)v\right].$ (12)
This shows that the adaptation of free-form flows for an invertible function $f$ to isometrically embedded manifolds is remarkably simple (see full proof in Section A.2; if the manifold is not isometrically embedded, add the corresponding term in Eq. 9):
@@
$\nabla_{\theta}\mathcal{L}_{\mathcal{M}\text{-NLL}}=\nabla_{\theta}\mathbb{E}_{x,v}\left[-\log p_{Z}(z)-v^{T}f_{\theta}^{\prime}(x)\texttt{SG}[{f_{\theta}^{-1}}^{\prime}(z)v]\right].$ (13)
@@
The only change is that $v$ must have covariance $RR^{T}$ rather than the identity. We achieve this by sampling standard normal vectors $\tilde{v}\in\mathbb{R}^{m}$ and then projecting them into the tangent space using the Jacobian of the projection function:
$v=\pi^{\prime}(f_{\theta}(x))\tilde{v}.$ (14)
Constructing $v$ like this fulfills the conditions of Theorem 3 because $\mathbb{E}_{v}[v]=0$, and:
$\operatorname{Cov}\left[v\right]=\mathbb{E}_{\tilde{v}}\left[\pi^{\prime}(f_{\theta}(x))\tilde{v}\tilde{v}^{T}\pi^{\prime}(f_{\theta}(x))^{T}\right]=\pi^{\prime}(f_{\theta}(x))\pi^{\prime}(f_{\theta}(x))^{T}=RR^{T}.$ (15)
Just like (Sorrenson et al., 2024; Draxler et al., 2024), we further normalize $v$ to reduce the variance of the trace estimator. Equation 13 now allows training invertible architectures on manifolds even if the volume change $\log|f_{\theta}^{\prime}(x)|$ is not tractable.
Free-form manifold-to-manifold neural networks
As discussed in Section 2, invertible architectures have to be specially constructed for each manifold. To overcome this limitation, we now soften the constraint that the learned model be analytically invertible. Instead, we learn a pair of free-form manifold-to-manifold neural networks, an encoder $f_{\theta}(x)$ and a decoder $g_{\phi}(z)$ as arbitrary functions on the manifold:
$f_{\theta}(x):\mathcal{M}\to\mathcal{M},\quad g_{\phi}(z):\mathcal{M}\to\mathcal{M}.$ (16)
We choose to fulfill Eq. 16 using feed-forward neural networks $\tilde{f}_{\theta},\tilde{g}_{\phi}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ working in an embedding space $\mathbb{R}^{m}$ of $\mathcal{M}$, but ensure that their outputs lie on the manifold by appending a projection $\pi:\mathbb{R}^{m}\to\mathcal{M}$, mapping points from the embedding space $\mathbb{R}^{m}$ to the manifold $\mathcal{M}$:
$f_{\theta}(x)=\pi(\tilde{f}_{\theta}(x)),\quad g_{\phi}(z)=\pi(\tilde{g}_{\phi}(z)).$ (17)
Figure 1 illustrates this for an example on a circle $\mathcal{M}=\mathbb{S}^{1}$.
Just like in the Euclidean case, we employ a reconstruction loss to make $f_{\theta}$ and $g_{\phi}$ approximately inverse to one another:
$\mathcal{L}_{\text{R}}=\mathbb{E}_{p_{\text{data}}}[\|g_{\phi}(f_{\theta}(x))-x\|^{2}].$ (18)
@@
We measure the distance in the embedding space; one can modify this to use an on-manifold distance (e.g. great circle distance for the sphere) but we find that ambient Euclidean distance works well in practice, since it is almost identical for small distances and this is the regime we work in.
This allows us to substitute ${f_{\theta}^{-1}}^{\prime}(z)\approx g_{\phi}^{\prime}(z)$ in Eq. 19:
$\nabla_{\theta}\mathcal{L}_{\mathcal{M}\text{-NLL}}\approx\nabla_{\theta}\mathbb{E}_{x,v}\left[-\log p_{Z}(z)-v^{T}f_{\theta}^{\prime}(x)\texttt{SG}[{g_{\phi}}^{\prime}(z)v]\right].$ (19)
Regularization and final loss
@@
$\mathcal{L}_{\text{U}}=\mathbb{E}_{x\sim\mathcal{U}(\mathcal{M})}[\|g_{\phi}(f_{\theta}(x))-x\|^{2}],$ (20)
helps ensure that we have a globally consistent mapping between the data and latent manifolds in low data regions. Secondly, the squared distance between the output of $\tilde{f}_{\theta}$ and its projection to the manifold:
$\mathcal{L}_{\text{P}}=\mathbb{E}_{p_{\text{data}}(x)}[\|\tilde{f}_{\theta}(x)-f_{\theta}(x)\|^{2}]$ (21)
discourages the output of $\tilde{f}_{\theta}$ from entering unprojectable regions, for example the origin when the manifold is $\mathbb{S}^{n}$. The same regularizations can be applied starting from the latent space.
@@
$\mathcal{L}=\mathcal{L}_{\mathcal{M}-\text{NLL}}+\beta_{\text{R}}\mathcal{L}_{\text{R}}+\beta_{\text{U}}\mathcal{L}_{\text{U}}+\beta_{\text{P}}\mathcal{L}_{\text{P}}$ (22)
where the gradient of $\mathcal{L}_{\mathcal{M}-\text{NLL}}$ is computed using Eq. 19, and $\beta_{\text{R}}$, $\beta_{\text{U}}$ and $\beta_{\text{P}}$ are hyperparameters. We give our choices in Appendix B.
@@ 5 Experiments
We now demonstrate the practical performance of manifold free-form flows on various manifolds. We choose established experiments to ensure comparability with previous methods, and find:
  • M-FFF matches or outperforms previous single-step methods. M-FFF uses a simple ResNet architecture, whereas previous methods were specialized to the given manifolds, hindering adoption to novel manifolds.
  • M-FFF is competitive with methods sampling in several steps at significantly faster inference speeds.
In our result tables, we mark as bold (a) the best method overall (both single- and multi-step), and (b) the best single-step method. We give all details necessary to reproduce the experiments in Appendix B. We run each experiment multiple times with different data splits and report the mean and standard deviation of those runs.
Synthetic distribution over rotation matrices
The group of 3D rotations $SO(3)$ can be represented by rotations matrices with positive determinant, i.e., all $Q\in\mathbb{R}^{3\times 3}$ with $Q^{T}Q=I$ and $\det Q=1$. We choose $\mathbb{R}^{3\times 3}$ as our embedding space and project to the manifold by solving the constrained Procrustes problem via SVD (Lawrence et al., 2019) (see Section B.2).
We evaluate M-FFF on synthetic mixture distributions proposed by De Bortoli et al. (2022) with $M$ mixture components for $M=16,32$ and $64$. Samples from one of the distributions and samples from our model are depicted in Fig. 3.
Table 3 shows that M-FFF outperforms the normalizing flow developed for $SO(3)$ by Liu et al. (2023), as well as the diffusion-based approaches for the mixtures $M=32$ and $64$.
Figure 3: Manifold free-form flows on a synthetic $SO(3)$ mixture distribution with $M=64$ mixture components proposed by De Bortoli et al. (2022). (Left) 10,000 samples each from the ground truth distribution and (right) our model. This visualization computes three Euler angles, which fully describe a rotation matrix, and then plot the first two angles on the projection of a sphere and the last by color (Murphy et al., 2021). We find that our model nicely samples from the distribution with few outliers between the modes.
Table 3: Test negative log-likelihood (NLL, $\downarrow$) on $SO(3)$ for multi-step and single-step methods. M-FFF consistently outperforms the specialized normalizing flow by Liu et al. (2023) on synthetic distributions of $SO(3)$ matrices, and outperforms multi-step methods in the cases with more mixture components. Multi-step baseline values are due to De Bortoli et al. (2022).
$M=16$ $M=32$ $M=64$ Fast inference?
Moser flow (Rozen et al., 2021) -0.85
@@
: 100 steps
$SO(3)$-NF (Liu et al., 2023) -0.81
$\pm$ 0.01
-0.12
$\pm$ 0.004
0.61
$\pm$ 0.01
M-FFF (ours) -0.87 $\pm$ 0.02 -0.21 $\pm$ 0.02 0.45 $\pm$ 0.02
Earth data on the sphere
Table 4: M-FFF significantly outperforms the previous single-step density estimator (Peel et al., 2001) on the sphere on real-world earth datasets in terms of negative log-likelihood (lower is better). Baseline values are collected from De Bortoli et al. (2022); Huang et al. (2022); Chen and Lipman (2024).
Volcano Earthquake Flood Fire Fast inference?
Riemannian CNF (Mathieu and Nickel, 2020) -6.05
$\pm$ 0.61
0.14
$\pm$ 0.23
1.11
$\pm$ 0.19
-0.80
$\pm$ 0.54
: $\sim$100 steps
Moser flow (Rozen et al., 2021) -4.21
$\pm$ 0.17
-0.16
$\pm$ 0.06
0.57
$\pm$ 0.10
-1.28
$\pm$ 0.05
: $\sim$100 steps
Stereographic score-based (De Bortoli et al., 2022) -3.80
$\pm$ 0.27
-0.19
$\pm$ 0.05
0.59
$\pm$ 0.07
-1.28
$\pm$ 0.12
: $\sim$100 steps
Riemannian score-based (De Bortoli et al., 2022) -4.92
$\pm$ 0.25
-0.19
$\pm$ 0.07
0.45
$\pm$ 0.17
-1.33
$\pm$ 0.06
: $\sim$100 steps
Riemannian diffusion (Huang et al., 2022) -6.61
$\pm$ 0.97
-0.40 $\pm$ 0.05 0.43
$\pm$ 0.07
-1.38
$\pm$ 0.05
: $>$100 steps
Riemannian flow matching (Chen and Lipman, 2024) -7.93 $\pm$1.67 -0.28
$\pm$ 0.08
0.42 $\pm$ 0.05 -1.86 $\pm$ 0.11 : 1000 steps
Mixture of Kent (Peel et al., 2001) -0.80
@@
M-FFF (ours) -2.25 $\pm$ 0.02 -0.23 $\pm$ 0.01 0.51 $\pm$ 0.01 -1.19 $\pm$ 0.03
Datset size 827 6120 4875 12809
We evaluate manifold free-form flows on spheres with datasets from the domain of earth sciences. We use four established datasets compiled by Mathieu and Nickel (2020) for density estimation on $\mathbb{S}^{2}$: Volcanic eruptions (NGDC/WDS, 2022b), earthquakes (NGDC/WDS, 2022a), floods (Brakenridge, 2017) and wildfires (EOSDIS, 2020).
Figure 1 shows an example for a model trained on flood data. As the reconstruction error sometimes does not drop to a satisfactory level we employ the method described in Section B.1 to ensure that the measured likelihoods are accurate. Table 4 shows that M-FFF again outperforms the specialized single-step model; the performance compared to multi-step methods is mixed. We think that multi-step models have an advantage on the considered data, as there are large regions of empty space between highly concentrated data points (see density and sample plots in Section B.3).
Torsion angles of molecules on tori
Table 5: M-FFF consistently outperforms normalizing flows specialized to tori (Rezende et al., 2020) on torus datasets, without requiring the development of a specialized architecture. In addition, our method comes close to the performance of the multi-step methods and even outperforms them on the Glycine dataset. Baseline values are due to Huang et al. (2022); Chen and Lipman (2024).
General Glycine Proline Pre-Pro RNA Fast inference?
Riemannian diffusion (Huang et al., 2022) 1.04
$\pm$ 0.012
1.97
$\pm$ 0.012
0.12
$\pm$ 0.011
1.24
$\pm$ 0.004
-3.70
$\pm$ 0.592
: $\sim$1000 steps
Riemannian flow matching (Chen and Lipman, 2024) 1.01 $\pm$ 0.025 1.90
$\pm$ 0.055
0.15
$\pm$ 0.027
1.18 $\pm$ 0.055 -5.20 $\pm$ 0.067 : 1000 steps
Mixture of power spherical (Huang et al., 2022) 1.15
@@
Circular Spline Coupling Flows (Rezende et al., 2020) 1.03 $\pm$ 0.01 1.91
$\pm$ 0.04
0.21
$\pm$ 0.08
@@
$\pm$ 0.04
-4.01
$\pm$ 0.24
M-FFF (ours) 1.03 $\pm$ 0.02 1.89 $\pm$ 0.05 0.17 $\pm$ 0.08 1.23 $\pm$ 0.04 -4.27 $\pm$ 0.09
@@
To benchmark manifold free-form flows on tori $\mathbb{T}^{n}$, we follow (Huang et al., 2022) and evaluate our model on two datasets from structural biology. We consider the torsion (dihedral) angles of the backbone of protein and RNA substructures respectively.
@@
We represent a tuple of angles $(\phi_{1},\dots,\phi_{n})\in[0,2\pi]^{n}$ by mapping each angle to a position on a circle: $X_{i}=(\cos\phi_{i},\sin\phi_{i})\in\mathbb{S}^{1}$. Then we stack all $X_{i}$ into a matrix $X\in\mathbb{R}^{n\times 2}$, compare Table 2.
@@
The first dataset is comprised of 500 proteins assembled by (Lovell et al., 2003) and is located on $\mathbb{T}^{2}$. The three dimensional arrangement of a protein backbone can be described by the so called Ramachandran angles (Ramachandran et al., 1963) $\Phi$ and $\Psi$, which represent the torsion of the protein backbone around the $N$-$C_{\alpha}$ and $C_{\alpha}$-$C$ bonds. The data is split into four distinct subsets General, Glycine, Proline and Pre-Proline, depending on the residue of each substructure.
@@
The second dataset is extracted from a subset of RNA structures introduced by Murray et al. (2003). As the RNA backbone structure can be characterized by seven torsion angles, in this case we are dealing with data on $\mathbb{T}^{7}$.
We report negative log-likelihoods in Table 5, finding that M-FFF outperforms a circular spline coupling flow, a normalizing flow particularly developed for data on tori (Rezende et al., 2020) as well as the multi-step methods on one of the datasets. In addition to the quantitative results, we show the log densities of the M-FFF models for the four protein datasets inFig. 5 in Section B.4 .
Toy distributions on hyperbolic space
We apply M-FFF to the Poincaré ball model, which embeds the 2-dimensional hyperbolic space $\mathbb{H}^{2}$ of constant negative curvature -1 in the 2-dimensional Euclidean space $\mathbb{R}^{2}$, as specified in Table 2. As this embedding is not isometric, and distances between points grow when moving away from the origin, we must include the last term of Eq. 9 when changing variables under a map on this embedded manifold.
We show that M-FFF can be applied to non-isometric embeddings using Eq. 9 and visualize learned densities in Fig. 1 and in Fig. 6 in Section B.5 for several toy datasets defined on the 2-dimensional Poincaré ball model. Further details can be found in Section B.5.
Manifold with non-trivial curvature
Finally, we follow Chen and Lipman (2024) and train M-FFF given by synthetic distributions on the Stanford bunny (Turk and Levoy, 1994) on the data provided with their paper, see Fig. 1. The natural embedding of this mesh is $\mathbb{R}^{3}$, and we train a separate neural network to project from the embedding space to the mesh. This ensures that the projection is continuously differentiable, which we identify to be important for stable gradients.
Table 6 shows that M-FFF performs well on this manifold, outperforming Riemannian flow matching in two out of three cases. This experiment underlines the flexibility of our model: We only need a projection function to the manifold in order to train a generative model.
Table 6: Test NLL on Stanford bunny data proposed by (Chen and Lipman, 2024), living on a manifold with nontrivial curvature (see Fig. 1). M-FFF outperforms the multi-step model for datasets with more modes.
$k=10$ $k=50$ $k=100$ Fast inference?
Riemannian Flow Matching (w/ diffusion) (Chen and Lipman, 2024) 1.16
$\pm$ 0.02
1.48
$\pm$ 0.01
1.53
$\pm$ 0.01
: 1000 steps
Riemannian Flow Matching (w/ biharmonic) (Chen and Lipman, 2024) 1.06 $\pm$ 0.05 1.55
$\pm$ 0.01
1.49
$\pm$ 0.01
: 1000 steps
M-FFF (ours) 1.21 $\pm$ 0.01 1.34 $\pm$ 0.01 1.28 $\pm$ 0.01